# One for All: Towards Training One Graph Model for All Classification Tasks

### v2

@@ 3.3 Graph prompting paradigm for graph in-context learning
To summarize, NOI represents the set of nodes related to the task, and the extracted NOI subgraph includes the neighborhood information of NOI. Then, the NOI prompt node summarizes information in the NOI by a graph learning model because all NOI nodes connect to the NOI prompt node. The NOI prompt node is later connected to a set of class nodes with text descriptions. After graph model processing, class node representations contain class information, task information, and NOI information, which can be used to make predictions independently, just like a prompted input to LLM contains the input, target, and task description.
We design a GNN capable of recognizing different edge types to serve as the graph model in OFA. The implementation details and training procedure of the model can be found in Appendix C.
Table 2: A comparison between OFA and related methods.
In-context Few-shot Zero-shot Cross-tasks Cross-domain GNN
LMMOL
GIMLET
InstructGLM
GraphText
NLGraph
GPT4GRAPH
ExpAsFeat
VNT
LLMForGraph
PRODIGY
GraphPrompt
All-in-One
OFA
@@ 4 Related works
The success of the LLM and prompt learning has enlightened many recent works that try to incorporate similar ideas to graphs. The first line of research tries to design prompt learning in the graph domain. Both VNT (Tan et al., 2023) and GraphPrompt (Liu et al., 2023c) introduce trainable prompt vectors to extract related information for different downstream tasks. GraphPrompt further utilizes the subgraph and link prediction to unify different task types. Prodigy (Huang et al., 2023) converts classification tasks to link prediction problems on prompt graphs, facilitating in-context graph learning. All-in-one (Sun et al., 2023) proposes to learn prompt graphs from downstream tasks. Our work also uses prompting to unify all classification tasks. More importantly, unlike existing work that still needs to train separate GNNs in different domains, we leverage language models to unify different tasks further and use one GNN to solve all tasks. The second line of research combines language models with GNNs. Some works directly apply LLMs to solve graph problems by describing the graph using natural language and feeding text description to LLMs, including InstructGLM (Ye et al., 2023), GraphText (Zhao et al., 2023c), NLGraph (Wang et al., 2023) and GPT4Graph (Guo et al., 2023). However, these methods also describe graph connections by texts, losing important structural features, while our method explicitly utilizes the information through GNNs. Very recently, GraphGPT (Tang et al., 2023) introduced a new method to encode the graph structure information with trainable vectors and an alignment module. It concatenates these trainable vectors along with the textual input to LLMs to improve the structure reasoning ability of LLMs. Some work for molecular graph classification proposes to only use SMILE molecule sequences to describe the molecular graph in input to the LLM, including LMMOL (Qian et al., 2023) and GIMLET (Zhao et al., 2023a). While they can perform in-context learning with LLMs, they are limited to the molecular graph domain. Other works use LLMs to encode the corresponding texts of nodes, such as paper abstracts, and apply GNN to the graph with encoded text embeddings, including ExpAsFeat (He et al., 2023) and LLMForGraph (Chen et al., 2023). GLEM (Zhao et al., 2023b) further introduces variational inference to improve the quality of test embedding during the training. These works focus on graph data that already have text features, whereas we are the first work that realizes that all graphs can be described by text and cross-domain data can be trained together in one model using LLMs. Table 2 summarizes the characteristics of different models. We can see that OFA is the most versatile model among existing works. More related works can be found in Appendix A.
@@ 5 Experiments
Cora Cora PubMed PubMed ogbn-arxiv Wiki-CS HIV
@@
Refer to caption Figure 3: Output embedding space of NOI prompt nodes on all datasets for OFA-joint-st.
Table 4: Results on supervised learning (second).
@@ 5.1 Cross-domain supervised learning
From the results, we have the following observations: (1) Both the independent and joint training achieve comparable or better results on all datasets compared to baseline methods. This indicates that using the text feature from the LLM encoder can obtain reasonable results for all graph datasets, including molecule graphs, which are generally not text-based in common sense. (2) OFA successfully enabled a single graph model to be effective on all graph datasets across different domains as the joint version with all different LLMs performs well on all datasets. Further, we can see that the joint version OFA-st achieves better results on most of the datasets compared to OFA-ind-st. This may indicate that by leveraging the text feature, the knowledge learned from one domain can be useful for the learning of other domains. Meanwhile, because of the proposed GPP, in-domain datasets with different tasks, such as molecular data, can be trained together, additionally boosting performance. (3) The comparison of different LLMs is interesting, generally speaking, a larger LLM can achieve better and more stable performance in joint training. We also observe a faster convergence for larger LLM (Llama2-13b). However, the margin is inapparent. Meanwhile, different LLMs seem specialized for different domains. For example, Llama2 achieves better performance than other LLMs in citation networks but e5-large-v2 achieve great result in molecular datasets. This may be attributed to different pertaining resources for different LLMs. We believe the effectiveness of LLMs requires further investigation and leave it to our future works. (4) Both the independent and the joint version of OFA achieve worse performance on Cora (node), Pubmed (node), and Wiki-CS compared to the baseline methods. This may be because the fixed-length vectors obtained from the frozen LLM encoder are less discriminative than one-hot encoding in some cases, which impedes the learning of the model. To further investigate the mechanism behind the OFA, we take the output embedding of NOI prompt nodes from OFA-joint-st for each dataset and project it to two-dimensional space. As shown in Figure 3, node embeddings from different domains are clearly separated. This demonstrates that the OFA model can represent data from different domains in different sub-spaces to process it. In Appendix 18, we conduct additional ablation studies to verify the effectiveness of the proposed GPP.
@@
To answer Q4, we design few-shot and zero-shot scenarios for node-level, link-level and graph-level tasks. We consider both transductive and transfer situations. In the transductive setting, given a dataset with label set $\mathcal{Y}$ where $\mathcal{Y}_{train}\cap\mathcal{Y}_{val}\cap\mathcal{Y}_{test}=\emptyset$, we use data associated with $\mathcal{Y}_{train}$ for training and evaluate on data associated with $\mathcal{Y}_{test}$. In the transfer setting, we examine the model’s capability to transfer knowledge to unseen datasets. We keep the joint training setup to train and evaluate a single model using all training datasets simultaneously. For simplicity, all experiments are conducted using OFA datasets generated from the sentence transformer. During training, we construct $N$-way $k$-shot tasks with varied $N\geq 2$ and $k\geq 0$, such that diverse few-shot and zero-shot tasks are involved for training. We include ogbn-arxiv, FB15K237, and Chemble as training sets, then evaluate two transductive settings: ogbn-arxiv and FB15K237, and four transfer settings: Cora, WN18RR, HIV, and PCBA. We present node/link/graph level results in Table 5 / 7 / 7, respectively. The model is denoted by OFA-joint-lr (low-resource).
@@
For node-level tasks, we compare with various baselines, including meta-learning models GPN (Ding et al., 2020), TENT (Wang et al., 2022c), and GLITTER (Wang et al., 2022b), as well as graph contrastive learning with fine-tuning proposed by TLP (Tan et al., 2022). For graph-level tasks, LLM-based models Galactica-1.3B (Taylor et al., 2022) and GIMLET (Zhao et al., 2023a) are considered. Note that while Prodigy Huang et al. (2023) is mentioned as the baseline for node and link level tasks, it follows a different setting, being trained on the MAG240M or Wiki datasets (Hu et al., 2021) for corresponding node or link tasks. OFA exhibits comparable or better performance than most existing works on few-shot tasks. Especially in the transfer setting of node tasks, where all baselines are trained and evaluated on the Cora dataset, OFA still shows comparable performance without any prior knowledge about the test dataset, illustrating its capability to generalize. Furthermore, our proposed GPP endows OFA with the ability to address zero-shot scenarios—a task generally impossible for most existing baseline models. OFA utilizes one single model to address all low-resource tasks across domains, demonstrating the ability of in-context learning.
@@
6 Limitations and future works
7 Conclusions
In this work, we propose OFA, the first solution towards building the foundation GNN model for learning on graphs. Specifically, the proposed text-attributed graphs, nodes-of-interest and graph prompt paradigm endow the OFA with the ability to perform various graph-related tasks from different domains with a single graph model. We evaluate OFA on supervised, few-shot, and zero-shot learning scenarios across different graph domains. The OFA show great results on most of the tasks and reveal great potential as the future foundation model on the graph.
@@ Appendix A Related Works (Extended)
Large Language Model and Prompt Learning. Since the ground-breaking advances in Large Language Models, including GPT (Brown et al., 2020; Wei et al., 2022) and LLaMa (Touvron et al., 2023), tremendous attention has been drawn to developing and using them. One particular approach is prompt learning. By prompting the LLMs according to a carefully designed paradigm, users can uncover LLM’s surprising capability in many difficult tasks. Chain-of-thoughts, by providing step-by-step reasoning examples, greatly improve LLMs’ reasoning power. Considerable efforts were also made to use prior semantic knowledge learned by LLMs to perform zero-shot and few-shot tasks (Reynolds & McDonell, 2021; Liu et al., 2023b; Chowdhery et al., 2022). The key advantage of prompting LLMs is that no further fine-tuning is required for the model to adapt to a new class of tasks. Innovated by this practical advantage, several works adapted the prompting idea to the GNN domain. VNT (Tan et al., 2023) offers a new method for few-shot node classification (FSNC) by integrating trainable virtual node embeddings into the original graph, which serves as a form of prompting. For each new FSNC task, the pre-trained graph transformer model can be frozen and only the embedding of virtual nodes needs to be trained. GraphPrompt (Liu et al., 2023c) also introduces trainable prompt vectors to extract related information for the downstream tasks. Meanwhile, it further utilizes the subgraph and link prediction to design a novel unsupervised pertaining method, which unifies the pertaining and downstream tasks. Prodigy Huang et al. (2023) converts classification tasks to link prediction problems on prompt graphs, facilitating in-context graph learning. All-in-one Sun et al. (2023) proposes to learn prompt graphs from data. Our work also uses prompting to unify all classification tasks. However, our works differ in two parts. First, unlike existing work that still needs to train separate GNNs in different domains, we leverage language models to unify different tasks further and use one GNN to solve all tasks. Second, most of the existing works cannot perform zero-shot learning as their prompt module needs to be re-trained for different downstream tasks. instead, OFA directly describes all tasks in a unified way, which allows the model to perform zero-shot learning without further training.
@@
The applications of GNN have also become prevalent. Much research focuses on the molecular property prediction domain and has achieved promising performance (Zhang et al., 2023a; Feng et al., 2023). GNNs have also become one of the primary tools in citation network analysis and mining (Kipf & Welling, 2017; Chien et al., 2022). Many efforts have also been made to adapt GNNs to the Knowledge Graph mining domain and achieved great success due to GNN’s efficiency and inductive learning capabilities (Zhu et al., 2021; Kong et al., 2022). While the promising results of these applications demonstrate the potential of GNNs, a critical problem is that they are tailored to the specified application. On the contrary, our work, by three carefully designed components, unifies all tasks into one GNN framework that allows large-scale training and inference across different domains. The proposed framework can be a foundation model in the graph learning domain.
@@
Language Models and Graph Neural Networks. The surge of foundational Large Language Models inspires several directions combining language models with GNNs. One of them directly applies LLMs to solve graph problems. They propose graph descriptive languages to represent the structural information as prompt texts and feed them to LLMs, like InstructGLM (Ye et al., 2023), GraphText (Zhao et al., 2023c), NLGraph (Wang et al., 2023), and GPT4Graph (Guo et al., 2023). Some works for molecular graph classification propose to only use SMILE molecule sequences to describe the molecular graph in input to the LLM, including LMMOL (Qian et al., 2023) and GIMLET (Zhao et al., 2023a). The exceptional power of LLMs enables these methods to perform difficult tasks such as zero/few-shot learning. However, such graph representation is implicit, and the LLM might not correctly capture the structural information but only summarize texts corresponding to the nodes that appear in the prompt. Such approaches will fall short in applications where graph structures are important. Very recently, GraphGPT (Tang et al., 2023) introduced a new method to encode the graph structure information with trainable vectors and an alignment module. It concatenates the trainable vectors along with textual input to LLms to improve the structure reasoning ability of LLMs. Another direction uses LLMs to encode the corresponding texts of nodes, such as paper abstracts, and apply GNN to the graph with encoded text embeddings (He et al., 2023; Chen et al., 2023). The high-quality text representation from LLMs allows these models to achieve better performance. Our approach adopts the same idea of using text embedding to represent graph entities (nodes and edges). However, we are the first work that drives the language model to full power by systematically unifying graph entities in different domains with the same language protocol and representing them in the same embedding space. This consequently grants cross-domain functionality that other GNN frameworks lack.
@@ B.1.1 Cora
Cora is a citation network contains papers and their citation relationship in computer science domain. The raw text data of the Cora dataset was collected from the GitHub repository provided in Chen et al. (2023). Each node in Cora represents a research paper from the computer science domain. The raw text feature of a node is the title and abstract of the respective paper. Every edge in the Cora dataset indicates the citation relationship between papers. Each node’s label corresponds to the category of the paper. Tasks that can be executed on Cora include predicting the paper’s category (node-level) or identifying missing citation links within the graph (link-level). Using the proposed processing protocol, we reformat all text features in Cora. Particularly, for each node category, we further use gpt-3.5-turbo(ChatGPT) to generate a description as additional information. In table 8, we show a processed example on Cora dataset.
@@
Node type
Text feature
Input graph node
feature node. literature category and description: $<$title$>$. $<$abstract$>$
Input graph edge
feature edge. co-citation
@@
Input graph edge
feature edge. citation
(OGBN-ARXIV/MAG240)
@@
Class node
prompt node. literature category and description: $<$category name$>$. $<$description$>$
Prompt node
prompt node. node classification of literature category.
@@
Class node
prompt node. two papers (do not have/have) co-citation
Prompt node
prompt node. link prediction on the papers that are cited together

### v3

@@ 3.3 Graph prompting paradigm for graph in-context learning
To summarize, NOI represents the set of nodes related to the task, and the extracted NOI subgraph includes the neighborhood information of NOI. Then, the NOI prompt node summarizes information in the NOI by a graph learning model because all NOI nodes connect to the NOI prompt node. The NOI prompt node is later connected to a set of class nodes with text descriptions. After graph model processing, class node representations contain class information, task information, and NOI information, which can be used to make predictions independently, just like a prompted input to LLM contains the input, target, and task description. The implementation details and training procedure of the model can be found in Appendix C.
@@ 4 Related works
The success of the LLM and prompt learning has enlightened many recent works that try to incorporate similar ideas to graphs. The first line of research tries to design prompt learning in the graph domain. Both VNT (Tan et al., 2023) and GraphPrompt (Liu et al., 2023c; Yu et al., 2023) introduce trainable prompt vectors to extract related information for different downstream tasks. HGPROMPT (Yu et al., 2024a) further extends GraphPrompt to heterogeneous graphs. Prodigy (Huang et al., 2023) converts classification tasks to link prediction problems on prompt graphs, facilitating in-context graph learning. All-in-one (Sun et al., 2023) proposes to learn prompt graphs from downstream tasks. Our work also uses prompting to unify all classification tasks. More importantly, unlike existing work that still needs to train separate GNNs in different domains, we leverage language models to unify different tasks further and use one GNN to solve all tasks. The second line of research combines language models with GNNs. Some works directly apply LLMs to solve graph problems by describing the graph using natural language and feeding text description to LLMs, including InstructGLM (Ye et al., 2023), GraphText (Zhao et al., 2023b), NLGraph (Wang et al., 2023) and GPT4Graph (Guo et al., 2023). However, these methods also describe graph connections by texts, losing important structural features, while our method explicitly utilizes the information through GNNs. Very recently, GraphGPT (Tang et al., 2023) introduced a new method to encode the graph structure information with trainable vectors and an alignment module. More related works can be found in Appendix A.
@@ 5 Experiments
Cora Cora [^1] The split we use differs from original one, see Apeendix D for details. PubMed PubMedfootnotemark: ogbn-arxivfootnotemark: Wiki-CS HIV
@@
Figure 3: Output embedding space of NOI prompt nodes on all datasets for OFA-joint-st.
@@ 5.1 Cross-domain supervised learning
From the results, we have the following observations: (1) Both the independent and joint training achieve comparable or better results on all datasets compared to baseline methods. (2) OFA successfully enabled a single graph model to be effective on all graph datasets across different domains as the joint version with all different LLMs performs well on all datasets. Further, we can see that the joint version OFA-st achieves better results on most of the datasets compared to OFA-ind-st. This may indicate that by leveraging the text feature and GPP, the knowledge learned from one domain/dataset can be useful for the learning of other domains/datasets. (3) The comparison of different LLMs is interesting, generally speaking, a larger LLM can achieve better and more stable performance in joint training. We also observe a faster convergence for larger LLM (Llama2-13b). However, the margin is less significant. Meanwhile, different LLMs seem specialized for different domains. For example, Llama2 achieves better performance in citation networks but e5-large-v2 achieves great results in molecular datasets. To further investigate the mechanism behind the OFA, we take the output embedding of NOI prompt nodes from OFA-joint-st for each dataset and project it to two-dimensional space. As shown in Figure 3, node embeddings from different domains are separated. This demonstrates that the OFA model can represent data from different domains in different sub-spaces to process it. In Appendix 18, we conduct additional ablation studies to verify the effectiveness of the proposed GPP.
@@
To answer Q4, we design few-shot and zero-shot scenarios for all levels of tasks. We consider both transductive and transfer situations. In the transductive setting, the task is to classify unseen classes on the same training graph. In the transfer setting, both the test graph and test classes are unseen. For simplicity, all experiments are conducted using OFA datasets generated from the sentence transformer. We train one few-shot model on various tasks varying $N$-way and $k$-shot, where $N\geq 2$ and $k\geq 0$. We include ogbn-arxiv, FB15K237, and Chemble as training sets, then evaluate two transductive settings: ogbn-arxiv and FB15K237, and four transfer settings: Cora, WN18RR, HIV, and PCBA. We present node/link/graph level results in Table 4 / 6 / 6, respectively. The model is denoted by OFA-joint-lr (low-resource).
@@
For node-level tasks, we compare with meta-learning models (Ding et al., 2020; Wang et al., 2022c; b) and graph contrastive learning model (Tan et al., 2022). For graph-level tasks, LLM-based models Galactica-1.3B (Taylor et al., 2022) and GIMLET (Zhao et al., 2023a) are considered. Prodigy Huang et al. (2023) follows a different setting, where the model is trained on the MAG240M or Wiki datasets (Hu et al., 2021) and transferred to the corresponding tasks. OFA exhibits comparable or better performance than most existing works on few-shot tasks. Especially in the transfer setting of node tasks, where all baselines are trained and evaluated on the Cora dataset, OFA still shows comparable performance without any prior knowledge about the test dataset, illustrating its capability to generalize. Furthermore, our proposed GPP endows OFA with the ability to address zero-shot scenarios—a task generally impossible for most existing baseline models. OFA utilizes one single model to address all low-resource tasks across domains, demonstrating the ability of in-context learning.
@@
6 Conclusions, limitations and future research
In this work, we propose OFA, the first solution towards building the foundation GNN model for learning on graphs. By showing great results on supervised, few-shot, and zero-shot scenarios, OFA reveals great potential as the future foundation model on the graph. Currently, OFA falls short of learning regression tasks and the cross-domain datasets are limited, we leave this to future work. More discussion can be found in Appendix F.
Acknowledgement
Hao Liu, Jiarui Feng, Lecheng Kong, and Yixin Chen are supported by NSF grant CBE-2225809. Muhan Zhang is supported by the National Key R&D Program of China (2022ZD0160303) and National Natural Science Foundation of China (62276003).
@@ Appendix A Related Works (Extended)
Large Language Model and Prompt Learning. Since the ground-breaking advances in Large Language Models, including GPT (Brown et al., 2020; Wei et al., 2022) and LLaMa (Touvron et al., 2023), tremendous attention has been drawn to developing and using them. One particular approach is prompt learning. By prompting the LLMs according to a carefully designed paradigm, users can uncover LLM’s surprising capability in many difficult tasks. Chain-of-thoughts, by providing step-by-step reasoning examples, greatly improve LLMs’ reasoning power. Considerable efforts were also made to use prior semantic knowledge learned by LLMs to perform zero-shot and few-shot tasks (Reynolds & McDonell, 2021; Liu et al., 2023b; Chowdhery et al., 2022). The key advantage of prompting LLMs is that no further fine-tuning is required for the model to adapt to a new class of tasks. Innovated by this practical advantage, several works adapted the prompting idea to the GNN domain. VNT (Tan et al., 2023) offers a new method for few-shot node classification (FSNC) by integrating trainable virtual node embeddings into the original graph, which serves as a form of prompting. For each new FSNC task, the pre-trained graph transformer model can be frozen and only the embedding of virtual nodes needs to be trained. GraphPrompt (Liu et al., 2023c; Yu et al., 2023) also introduces trainable prompt vectors to extract related information for the downstream tasks. Meanwhile, it further utilizes the subgraph and link prediction to design a novel unsupervised pertaining method, which unifies the pertaining and downstream tasks. MultiGPrompt Yu et al. (2024b) proposes a multi-task pre-training and prompting framework, effectively harnessing diverse pretext task knowledge for enhanced graph-based analysis. Prodigy Huang et al. (2023) converts classification tasks to link prediction problems on prompt graphs, facilitating in-context graph learning. All-in-one Sun et al. (2023) proposes to learn prompt graphs from data. Our work also uses prompting to unify all classification tasks. However, our works differ in two parts. First, unlike existing work that still needs to train separate GNNs in different domains, we leverage language models to unify different tasks further and use one GNN to solve all tasks. Second, most of the existing works cannot perform zero-shot learning as their prompt module needs to be re-trained for different downstream tasks. instead, OFA directly describes all tasks in a unified way, which allows the model to perform zero-shot learning without further training.
@@
The applications of GNN have also become prevalent. Much research focuses on the molecular property prediction domain and has achieved promising performance (Zhang et al., 2023a; Feng et al., 2023; Kong et al., 2023; Feng et al., 2022). GNNs have also become one of the primary tools in citation network analysis and mining (Kipf & Welling, 2017; Chien et al., 2022). Many efforts have also been made to adapt GNNs to the Knowledge Graph mining domain and achieved great success due to GNN’s efficiency and inductive learning capabilities (Zhu et al., 2021; Kong et al., 2022). While the promising results of these applications demonstrate the potential of GNNs, a critical problem is that they are tailored to the specified application. On the contrary, our work, by three carefully designed components, unifies all tasks into one GNN framework that allows large-scale training and inference across different domains. The proposed framework can be a foundation model in the graph learning domain.
@@
Language Models and Graph Neural Networks. The surge of foundational Large Language Models inspires several directions combining language models with GNNs. One of them directly applies LLMs to solve graph problems. They propose graph descriptive languages to represent the structural information as prompt texts and feed them to LLMs, like InstructGLM (Ye et al., 2023), GraphText (Zhao et al., 2023b), NLGraph (Wang et al., 2023), and GPT4Graph (Guo et al., 2023). Some works for molecular graph classification propose to only use SMILE molecule sequences to describe the molecular graph in input to the LLM, including LMMOL (Qian et al., 2023) and GIMLET (Zhao et al., 2023a). The exceptional power of LLMs enables these methods to perform difficult tasks such as zero/few-shot learning. However, such graph representation is implicit, and the LLM might not correctly capture the structural information but only summarize texts corresponding to the nodes that appear in the prompt. Such approaches will fall short in applications where graph structures are important.
Table 7: A comparison between OFA and related methods.
In-context Few-shot Zero-shot Cross-tasks Cross-domain GNN
LMMOL
GIMLET
InstructGLM
GraphText
NLGraph
GPT4GRAPH
ExpAsFeat
VNT
LLMForGraph
PRODIGY
GraphPrompt
All-in-One
OFA
Very recently, GraphGPT (Tang et al., 2023) introduced a new method to encode the graph structure information with trainable vectors and an alignment module. It concatenates the trainable vectors along with textual input to LLms to improve the structure reasoning ability of LLMs. Another direction uses LLMs to encode the corresponding texts of nodes, such as paper abstracts, and apply GNN to the graph with encoded text embeddings (He et al., 2023; Chen et al., 2023). The high-quality text representation from LLMs allows these models to achieve better performance. Our approach adopts the same idea of using text embedding to represent graph entities (nodes and edges). However, we are the first work that drives the language model to full power by systematically unifying graph entities in different domains with the same language protocol and representing them in the same embedding space. This consequently grants cross-domain functionality that other GNN frameworks lack. Table 7 summarizes the characteristics of different models. We can see that OFA is the most versatile model among existing works.
@@ B.1.1 Cora
Cora is a citation network that contains papers and their citation relationship in the computer science domain. The raw text data of the Cora dataset was collected from the GitHub repository provided in Chen et al. (2023). Each node in Cora represents a research paper from the computer science domain. The raw text feature of a node is the title and abstract of the respective paper. Every edge in the Cora dataset indicates the citation relationship between papers. Each node’s label corresponds to the category of the paper. Tasks that can be executed on Cora include predicting the paper’s category (node-level) or identifying missing citation links within the graph (link-level). Using the proposed processing protocol, we reformat all text features in Cora. Particularly, for each node category, we further use gpt-3.5-turbo(ChatGPT) to generate a description as additional information. In table 8, we show a processed example on Cora dataset.
@@
Node type Text feature
Input graph node feature node. literature category and description: $<$title$>$. $<$abstract$>$
Input graph edge feature edge. co-citation
@@
Input graph edge feature edge. citation
(OGBN-ARXIV)
@@
Class node prompt node. literature category and description: $<$category name$>$. $<$description$>$
Prompt node prompt node. node classification of literature category.
@@
Class node prompt node. two papers (do not have/have) co-citation
Prompt node prompt node. link prediction on the papers that are cited together